Mar_11_2004_L1694.txt Shift L1694 Start at 2004:071 12:00 GMT = 2004 Mar 11 7:00 am EDT End 2004 Mar 11 10:15 am EDT SOT shift report 2004-03-011 L1694 http://asc.harvard.edu/acis/SOT_reports/sot-shift.html or /proj/web-cxc/htdocs/acis/SOT_reports/Mar_11_2004_L1694.txt SOT Lead Bill Podgorski SOT SC/ACA FOT-SI HRC-SI ACIS-SI OTHER Plan: - A 24 hour turn-around TOO was proposed and accepted yesterday shortly after noon. The TOO, '1ES 1028+511' (obsid 5036) is a blazar in outburst, observed with HRC-S/LETG. A set of TOO loads, MAR1104A, was produced and approved. The loads were uplinked in this morning's pass, replacing the MAR0804C loads. The TOO loads will start at about 071/1322, after obsid 4694. Soft proton levels are low (under 100) but there is a lot of backscatter noise. 3-day Solar-Geophysical Forecast issued Mar 10 at 22:00 UTC Solar Activity Forecast: Solar activity is expected to be low. Region 570 may produce occassional C-class flares. Geophysical Activity Forecast: Geomagnetic activity is expected to be at unsettled to active conditions, with isolated minor storm conditions possible on 11 March and early into 12 March. Activity should decrease to quiet and unsettled levels on 13 March as the coronal hole rotates out of geoeffective position. Seq # ObsID It Target start time SI Grat RA dec roll Pitch ------------------------------------------------------------------------------------------------------------ MAR0804C: 600349 P1 04694 0 NGC 3521 2004:071:10:36:42.103 10.0 ACIS-S NONE 166.4335 -0.0329 303.49 173.51 600347 P1 04692 0 NGC 3384 2004:071:13:56:01.646 10.0 ACIS-S NONE 162.0685 12.6505 227.75 166.57 MAR1104A: TOO load 900343 05036 0 1ES 1028+511 2004:071:14:23:54.772 150.0 HRC-S LETG 157.8476 50.9144 198.86 131.09 ---- T_E35 CAL-ER (60340) 2004:073:08:41:55.453 8.0 -- -- 176.5000 25.4000 172.61 157.19 ---- GG_35 CAL-ER (60339) 2004:073:14:18:36.465 3.6 -- -- 162.0000 22.0000 212.39 157.49 ---- T_X3600 CAL-ER (60338) 2004:073:19:48:24.658 19.8 -- -- 162.0000 22.0000 212.71 157.32 600387 05309 1 M101 2004:074:02:05:34.154 71.5 ACIS-S NONE 210.7909 54.3750 141.01 120.20 590364 05166 1 G21.5-09[S3,-120,-1. 2004:074:22:32:10.880 10.0 ACIS-S NONE 278.4055 -10.5711 90.29 76.53 ORBITAL EVENTS: 2004:071:02:51:19.417 634 EPERIGEE ORBIT PERIGEE 2004:072:10:34:39.954 635 EAPOGEE ORBIT APOGEE 2004:073:18:18:31.783 635 EPERIGEE ORBIT PERIGEE EVENTS DONE OBSID -------------- ---- ----- Start Shift 071:1200 AOS DSS-16 (Goldstone 26m) 071:1245 Status at AOS was as expected 4694 (see note 1) CAP 913 ran to replace MAR0804C with TOO loads 071:1255 Swap to DSS-27 due to problems 071:1430 LOS DSS-27 071:1515 End of shift 071:1515 NOTES/ISSUES/PROBLEMS --------------------- 1) Status at AOS: UTC 2004:071:12:51:04 (Mar11) f_ACE 2.47e+01 F_CRM 9.02e+07 Kp 4.0 R km 80921A OBT 2004:071:12:50:45 CTUVCDU 13641983 OBC s/w NRML FMT2_NORM CPEstat NORM OBT 195396645.20 ONLVCDU 13641968 OBC Errs 0 OBSID 4694 EPState SUN SIM TTpos 75623 HETG Angle 79.46 PCADMODE NPNT RA 166.432 Bus V 29.61 SIM FApos -467 LETG Angle 77.59 PCONTROL NPNT Dec -0.033 Bus I 28.24 AOFSTAR GUID Roll 303.490 ACA Object FFFSSSSS Dither ENAB HRC-I HV ON ACA ImFunc TTTTTTTT Dith Yang -3.68 Yaw Rate -0.19 HRC-S HV OFF ACA CCD Temp -15.6 Dith Zang 4.64 Pitch Rate -0.06 OBSMode NEXT ACA Int Time 1.698 Roll Rate -0.02 Door OP UNAC AOACSTAT OK FSS SunBeta NSUN Door CL UNAC FSS Alfa -49.62 Batt 1 SOC 100.00% Avg HRMA Temp 70.61 FSS Beta -49.72 Batt 2 SOC 100.00% ACIS Stat7-0 64 Avg OBA Temp 51.29 SA Resolv 173.61 Batt 3 SOC 100.00% Cold Rad -127.9 OBA Tavg Fault NFLT SA Sun Pres ILLM Warm Rad -82.0 OBA Trng Fault NFLT +Y SA Amps 13.72 RadMon ENAB HRMA power 9.60 SCS 128 ACT -Y SA Amps 14.30 EPHIN Geom LARG OBA power 131.50 SCS 129 INAC +Y SA Temp 115.12 E150 1.0 SCS 130 ACT -Y SA Temp 115.12 E300 0.0 Roll Mom. 1.955 SCS 107 INAC E1300 0.0 Pitch Mom. 9.862 UpL Cmd Acc 20609 EPH A-Leak 1.0200 P4GM 0.0 Yaw Mom. -14.962 Cmd Rej A 15 EPH B-Leak 0.4200 P41GM 0.0 EPH temp 68.08 Gyro 1 Curr 1 9.60 Roll Bias -0.6170 Gyro 1 Curr 2 9.60 Pitch Bias -1.8939 Gyro 2 Curr 1 103.20 Yaw Bias -0.9423 Gyro 2 Curr 2 100.80 Explanation Current Data: UTC 2004:071:12:51:04 (Mar11) f_ACE 2.47e+01 F_CRM 9.02e+07 Kp 4.0 R km 80921A Currently scheduled FPSI, OTG : ACIS-S NONE Estimated Kp : 4.00 ACE EPAM P3 Proton Flux (p/cm^2-s-sr-MeV) : 2.47e+01 GOES-10 P2 flux, in RADMON P4GM units : 0.00 GOES-12 P2 flux, in RADMON P4GM units : 0.00 GOES-10 P5 flux, in RADMON P41GM units : 0.01 GOES-12 P5 flux, in RADMON P41GM units : 0.01 Orbit Start Time : 2004:071:02:54:00 Geocentric Distance (km), Orbit Leg : 80765 A CRM Region : 3 (Magnetosphere) External Proton Flux (p/cm^2-s-sr-MeV) : 8.82519000000e+02 Attenuated Proton Flux (p/cm^2-s-sr-MeV) : 8.82519000000e+02 External Proton Orbital Fluence (p/cm^2-sr-MeV) : 1.97392628580e+10 Attenuated Proton Orbital Fluence (p/cm^2-sr-MeV) : 9.01614396000e+07 HRC IM PS = ON Stp 42/53 SP PS = OFF Shield 1 HV = OFF Shield 2 HV = ON Stp 5 Valid Event Rate 1 = 0 Valid Event Rate 2 = 0 Shield Event Rates = 3100 ACIS FP Temp = -119.7 DEA Pwr = 49.3 DPA Pwr = 40.3 Science State = Science Active 2) Notes from 9:00am meeting: Operations (Ken): - 1 supports over last 24 hours, 1 34m, 1 26m - 3 LS uplinked, (TOO loads) active thru 075/0118 - Issue with DSS-46 pass yesterday afternoon, similar to ond on day 069. - Dumps are now current Mission Planning (Ken): - See above discussion of TOO load. - MAR1504 loads being reworked to include the first part of the M101 observation (obsid 4732). - Products out later today S/C Engineering(Dan): - All OK on S/C - CAP review for momentum dump, maybe later today (or tomorrow) SOT(Bill): - See above notes on radiation. DSOPS(Jen): - Several QL's (see below) GOT(Mike): - All OK IST(Chris): - All OK Meetings/events: - 9:15am, IST Tagup - 10:00am, Linux migration meeting - 10:00am, SMS CCB - 1:00pm, Contractor telecon - 2:30pm, FDB CANCELLED this week - 4:00pm, OFLS MP meeting Roger: - Nice job on TOO - MIT visitor this afternoon 3) Quick look data: Seq # ObsID It Target start time SI Grat Comment ------------------------------------------------------------------------------------------- MAR0804C: 200278 04509 0 Spica (Alpha Vir) 2004:069:02:17:22.538 90.0 ACIS-S LETG 200288 P2 05180 0 HR 7012 2004:070:03:50:06.162 3.0 ACIS-I NONE 600404 04749 1 NGC 1313 X-2 Obs 1 2004:070:05:08:06.162 5.0 ACIS-S NONE 700952 P2 04870 0 2QZ J031056.9-305901 2004:070:06:59:26.162 15.0 ACIS-S NONE 800437 P2 04994 0 RXCJ0528.9-3927 2004:070:11:37:26.162 22.6 ACIS-I NONE